# Graph Reinforcement Learning in Power Grids: A Survey

### v1

@@ 5.1.1 Operation Control
Most approaches evaluate their performance on IEEE grids with a size ranging from 5 to 300 buses. The authors evaluate the presented approaches with metrics like voltage deviation, network energy loss, or voltage violation rates. Comparisons typically include traditional optimization methods and other deep RL approaches as benchmarks.
@@
The advantage of GNNs over dense-based RL agents is demonstrated in many of the approaches analysed. Starting with Lee et al., 2022b , the authors tested their method on PowerGym Fan et al., (2022) grids with 13, 34, 123, and 8500 nodes. Comparing the performances, their GNN-based PPO method slightly outperformed the dense PPO on smaller grids but converged slower on larger ones due to message-passing nature of the GNN layers. Additionally, experiments with missing or noisy observations showed that the superiority of the GNN-based approach in terms of robustness. Moreover, the paper finds voltage regulators impact the grid globally, while batteries and capacitors have local effects. To address this, they add edges between nodes with voltage regulators and use a local readout function for the controllable nodes, improving the robustness and performance of the GNN-based PPO approach. Similar to Lee et al., 2022b , Wang et al., (2023) compared their graph-PPO with a dense-PPO method on IEEE grids of 33 to 123 buses, demonstrating near-optimal performance and a better scalability.
@@
Similar grid sizes are addressed by Li et al., (2023), who significantly outperform optimisation-based benchmarks. Their GRL approach achieves faster inference times, higher rewards, lower voltage fluctuation rates, and higher renewable accommodation. Furthermore, is has a faster convergence due to the spatio-temporal attention in their GNN architecture. Their analysis of the attention masks shows strong connections to high-power buses and that 10 time steps are sufficient to capture necessary temporal information, as correlation weakens beyond this.
@@
Yan et al., (2023) compare their primal-dual GRL model with single-agent and multi-agent DDPG methods and their primal-dual variants. The GRL method effectively minimizes network energy losses and voltage deviations, even with noisy or incomplete data. Traditional volt-var optimization, while performing best with its access to global information, struggles with noisy or missing data. Similarly, Mu et al., (2023) find their GNN-based approach achieves more stable voltages and fewer violations compared to optimization methods and multi-agent DDPG on 33- and 141-bus grids. Further, it showed robustness to line and bus deletions. Wu et al., 2023a also show improved voltage profiles and reduced power losses on modified IEEE 33-node and 25-node systems compared to dense-based, CNN-based DDPG, and conventional optimization methods.
Wu et al., 2022c focused on different scenarios of outage attacks in 33- and 119-node systems with one-year smart meter readings and PV generation of a real distribution system. They showed that their approach mitigates oscillations from cyber-attacks even when 50% of inverters were corrupted, though they did not benchmark against other methods. Therefore, these results are difficult to interpret. Particularly, the utlilization of GNNs remains to be validated.
Xu et al., 2022b evaluated their results only against algorithmic solutions, showing their method was faster than the heuristics and close to optimal, with improvements from using GCN, branch exchange, and action separation. In an ablation study, they evaluate the benefits of the individual components of the proposed framework.
@@
Finally, Cao et al., (2023) conducted comparative tests on the IEEE 33-node system, evaluating their physics-informed GAT-SAC approach against several control methods, including SAC variants and GCN-SAC. Their method outperformed others in reducing voltage deviations and maintaining safe voltage levels, especially under noisy conditions. Ablation studies emphasized the importance of the GAT-based network and the added robustness from the deep autoencoder. Tests on the IEEE 119-node system confirmed the method’s scalability and effectiveness in larger networks.
5.1.2 Distribution Grid Control in Emergency Situations
@@
Another way to control grid voltages is load shedding, which involves deliberately disconnecting certain loads. This drastic measure is usually a last resort to avoid total blackouts. Conversely, in case of a full or partial blackout occurs, system restoration is needed to restart the grid. This involves coordinated steps to reconnect loads, restore generation capacity, and ensure the integrity of transmission and distribution networks. In the following, we present three load shedding and one system restoration approach that utilize GRL.
@@
The goal is to develop a GRL agent that learns to make load shedding decisions, as proposed by Hossain et al., (2021), Zhang et al., (2023), and Pei et al., (2023), which present three different GRL approaches for handling critical system loads.
@@
The problem is modeled as an MDP, predicting actions based on grid state observations like power demand, generation, voltage measurements, and topology. Pei et al., (2023) further include historical node voltages. As for the decision-making, Hossain et al., (2021) and Zhang et al., (2023) model decisions as binary actions (load shedding or not), while Pei et al., (2023) considers shedding 5% or 10% of the load. All three approaches focus on heavy load nodes, with action space scaling accordingly. Zhang et al., (2023) also includes switching lines on and off. The reward in all approaches aims to maintain stable voltage levels and prevent system collapse. For instance, Hossain et al., (2021) gives a high negative reward if voltage stays below 95% of the nominal after 4 seconds, and a positive reward for voltages within predefined levels. The reward also minimizes load shedding or maximizes power supply (Zhang et al., (2023)) and penalizes actions violating system constraints.
@@
RL algorithms used include DDQN with an $\epsilon$-greedy strategy and experience replay (Hossain et al., (2021) and Pei et al., (2023)), with Pei et al., (2023) also using a dueling architecture. Zhang et al., (2023) uses PPO with a hybrid policy network combining a fully connected network and a GCN.
@@
In all papers, the distribution grid is modeled as a graph, with buses as nodes. This includes elements like substations, loads, and generators that connected by edges representing power lines and transformers. Node features include grid state data such as voltage measurements and loads. All approaches use a GCN for feature extraction, but with different architectures. Hossain et al., (2021) uses a simple GCN with three graph convolution layers, followed by a fully connected layer. Zhang et al., (2023) employs a graph capsule network (cf. Sec. 3) that embeds whole graphs and stores additional neighborhood statistics to avoid information loss. A node embedding then contains higher-order statistics for each feature within the local neighbourhood as well as compressed node labels, similar to Weisfeiler-Lehman colourings Verma and Zhang, (2018). Conversely, Pei et al., (2023) uses GraphSAGE, which samples and aggregates the local neighborhood of nodes.
The presented papers conduct experiments on IEEE case example grids. Hossain et al., (2021) and Pei et al., (2023) use the 39-bus system, while Zhang et al., (2023) tests on the modified IEEE 13-bus and 34-bus systems. Each method trains on different topological configurations with random fault locations.
@@
All three approaches outperform feed-forward neural networks in convergence and average reward, especially on unseen topologies. Hossain et al., (2021) evaluates on 32 topologies. Pei et al., (2023) also tests their GNN-based method on the 300-bus system, managing larger action spaces better than feed-forward networks. Additionally, Zhang et al., (2023) and Pei et al., (2023) benchmark against traditional optimization techniques, achieving near-optimal performance with faster computation. Pei et al., (2023) demonstrates superior adaptability and efficiency, favoring GraphSAGE over a classical GCN. These methods enhance voltage control under varying grid topologies, showing the robustness and efficiency of GNNs compared to conventional methods.
@@
On a different note, in case earlier mitigation actions fail and a blackout occurs, rapid system restoration is crucial to reconnect loads and restart the grid promptly. Zhao and Wang, (2022) approach this optimization with a routing model, treating generators as energy depots to connect to as many loads as possible. They employ a multi-agent approach where a Q-Network guides actions, using an encoder to observe generator capacities, switch statuses, and load conditions. A GCN extracts features from local and neighboring agents for the Q-Networks. This method outperforms single-agent DQN and other multi-agent baselines using feedforward networks, as well as CPLEX, a mathematical optimization method, in terms of accuracy and speed. Case studies validate this approach on IEEE 123 and 8500 node test systems Zhao and Wang, (2022).
@@ 5.2 Other Use Cases
Two additional use cases for distribution grids include loss minimization and economic dispatch. Loss minimization focuses on optimizing the power grid state by adjusting topology or generator set points to reduce losses in branch elements. Economic dispatch, on the other hand, aims to minimize the operational costs of generators.
Jacob et al., (2022) focuses on Distribution Network Reconfiguration (DNR) to minimize losses and enhance resilience by dynamically reconfiguring the network’s topology through sectionalizing and tie line switches. This adapts the distribution network to changing conditions like demand fluctuations or disruptions. Their states include the current network topology, status of connections (open or closed), power demands at buses, generation outputs, bus voltages, branch currents, and total network loss for efficiency assessment. Rewards in Jacob et al., (2022) incentivize loss reduction and compliance with topological constraints, penalizing actions that disconnect loads or alter the radial structure, while rewarding feasible network exploration
Meanwhile, Chen et al., (2023) optimizes economic dispatch in power systems with high renewable energy penetration for economic efficiency and stability under dynamic and uncertain conditions, involving adjustments in generation, renewable output, and ESS charging/discharging power. The states include load demands, generation and RES outputs, ESS state of charge, and the current timestep, offering a comprehensive view of the power grid’s dynamic state. Rewards are designed to optimize economic dispatch and system stability, combining economic operation costs with a stability component addressing voltage violations, encouraging actions that balance efficiency and stability.
Regarding the graph representation, the nodes in Jacob et al., (2022) represent substations and buses connected by edges for power lines and transformers. They employ a Capsule-based Graph Convolutional Network (GCAPCN) (cf. Sec. 3) to capture both local and global graph features. Likewise, Chen et al., (2023) adopts a power system graph representation where nodes represent buses and edges represent power system branches. They utilize a GCN to process node features including load demand, generation outputs, RES outputs, ESS state, and timestep information, mapping the system state to an action policy through layers of transformation and feature extraction.
Table 3: Overview of other relevant GRL approaches proposed for distribution grids. The RL algorithm, the action as well as the GNN type are specified. The focus defines the objective of the proposed approaches.
Model RL Action GNN Focus
@@
For loss minimization, Jacob et al., (2022) use the PPO algorithm for training. The GCAPCN encodes grid details such as losses, bus voltages, branch currents, and topology for the policy network, while a feedforward neural network serves the value network. In contrast, Chen et al., (2023) apply the SAC algorithm with the GCN layers in both the actor and critic roles to optimize dispatch policies off-policy.
The approaches are evaluated as follows: Jacob et al., (2022) compare their GCAPS-RL method against two conventional baseline methods, Mixed-integer Second-order Conic Programming (MISOCP) and binary particle swarm optimization (BPSO), as well as a feed-forward approach to assess the benefits of GNNs. Experiments are conducted on modified IEEE distribution test networks (13-bus and 34-bus systems) using power flow simulations. GCAPS-RL demonstrates superior real-time decision-making and adherence to topological constraints, outperforming the feed-forward counterpart particularly in the 34-bus system.
Meanwhile, Chen et al., (2023) perform case studies on a modified IEEE39 power system configuration with conventional generators, renewable sources, and an ESS. Their GRL approach is compared against Optimal Solution with Perfect Information (OSPI) and Model Predictive Control (MPC) policies with varying prediction horizons, as well as feed-forward SAC. GRL exhibits strong convergence, effective policy performance, and superior scalability when the system topology changes. It achieves near-optimal results and reduces costs compared to all baselines, with efficient computation meeting practical requirements, especially under extended prediction horizons.
@@ 5.3 Discussion
Voltage control and grid control in emergency situations using DRL techniques encompass a variety of approaches and considerations. One key aspect is the design of reward functions, which typically address voltage deviations but may also incorporate additional factors such as the integration of renewables or power loss. However, balancing these objectives poses a non-trivial challenge as already mentioned in Sec. 4.5.
In terms of control devices, most methods focus on PV inverters or other actuators, though some also consider ESS or adjust the active power of generators. Notably, Xu et al., 2022b stands out for its consideration of topological actions, offering a unique perspective in this domain.
The control strategies for distribution grids that are divided into zones or microgrids involve multi-agent setups, where each agent corresponds to a distinct zone. Here, the treatment of global system knowledge varies, with approaches either training globally and evaluating locally or utilizing a centralized critic for action evaluation. In this setup GNNs can be used to fuse the information from different agents as presented in Mu et al., (2023).
The implementation of GNNs varies widely across studies, with no consensus on whether to implement them in the actor, critic, or both. To map node embeddings to actionable vectors, facilitating dimensionality reduction and action prediction, the readout methods vary, ranging from fully connected neural networks to autoencoders to 1D convolutions. Weight sharing between surrogate models and actors or critics is observed in some approaches. While graph representations remain largely consistent, GNN architectures exhibit significant diversity. Spectral GCNs are prevalent, although less common in other domains. This might be due to the fact that spectral convolutions come from signal processing and thus are closer to electrical engineering. Some approaches also utilize temporal information by including historical data, but most approaches consider static information only. Multiple studies introduce customized Graph Structure Operatorss which make use of domain knowledge about certain properties of the system. The respective approaches show that the feature extraction benefits from the incorporation of known system dynamics. Domain knowledge can also be used for the augmentation of the graph representation of grids to define edge weights and determine critical connections in beforehand.
Also already pointed out in Sec. 4.5, GNNs have outperformed fully connected neural networks in terms of transferability and adaptability to topology changes. Experiments with deleted grid elements or completely different grid structures can be handled without a drastic drop in performance, which is not the case for fully connected neural networks.
The ability of GNNs to handle noisy or missing data effectively is frequently cited as a motivation for their use. In the according experiments including generator failures, deleted lines or nodes, all works demonstrate the robustness of GNNs. In real grid operation, faulty sensor data or unknown parameters are the normal case rather than an exception, therefore this is a clear advantage.
@@
The experiments and evaluations of these approaches encompass a wide range of considerations. Most studies conduct experiments on IEEE grids, with grid sizes varying from small-scale systems with as few as 5 buses to large-scale networks with up to 8500 nodes. The grid sizes exhibit considerable variation, impacting the scalability and generalization of approaches. Notably, some methods demonstrate effective performance even without access to global information, underscoring the robustness of certain DRL strategies in this context. Evaluation metrics typically include voltage deviation, network energy loss, or voltage violation rates, reflecting the overarching objective of voltage control and load shedding. Traditional optimization techniques and heuristics serve as benchmarks in many studies, highlighting the comparative performance and efficiency gains of DRL-based approaches. Additionally, comparisons with other DRL methods, including dense-based and CNN-based approaches, shed light on the relative advantages of graph-based methodologies. The experiments often demonstrate the scalability, stability, and real-time deployment capabilities of the proposed frameworks, showcasing their effectiveness in addressing complex power system control challenges.

### v3

@@ 5.1.1 Operation Control
Most approaches evaluate their performance on IEEE grids ranging in size from 5 to 300 buses. The data for the injections can be randomly sampled or correspond to historical time series of power generation, mostly from photovoltaic. The authors evaluate the presented approaches using metrics such as voltage deviation, network energy loss, or voltage violation rates. Comparisons typically include traditional optimization methods, heuristics, and other deep RL approaches as benchmarks.
@@
The advantage of GNNs over dense-based RL agents is evident in several studies. Lee et al., 2022b tested their GNN-based PPO on PowerGym grids ranging from 13 to 8500 nodes, showing better performance and robustness, especially with noisy and missing data. In addition, the paper finds that voltage regulators affect the grid globally, while batteries and capacitors have local effects. To address this, the authors add edges between nodes with voltage regulators and use a local readout function for the controllable nodes, improving the robustness and performance of the GNN-based PPO approach. Similarly, Wang et al., (2023) compared their graph PPO with a dense PPO method on IEEE grids from 33 to 123 buses, demonstrating near-optimal performance and better scalability.
@@
Similar grid sizes are addressed by Li et al., (2023), whose GRL approach significantly outperforms optimization-based benchmarks with faster inference, higher rewards, lower voltage fluctuations, and greater renewable energy accommodation. Their spatio-temporal attention exhibits a faster convergence with the attention masks, indicating strong connections to high-power buses.
@@
Yan et al., (2023) found that their primal-dual GRL model minimized energy losses and voltage deviations and outperformed single-agent and multi-agent DDPG methods, especially with noisy data. Similarly, Mu et al., (2023) reported robustness to line and bus deletions as well as stable voltages and fewer violations on 33-bus and 141-bus grids when compared to optimization methods and multi-agent DDPG. This is the same for Wu et al., 2023a , whose approach improved voltage profiles and reduced power losses on IEEE 33-node and 25-node systems compared to dense-based and CNN-based DDPG and conventional optimization.
Wu et al., 2022c mitigated oscillations from cyber-attacks in 33- and 119-node systems, showing effective mitigation even with 50% of inverters compromised. They did not benchmark against other methods, hence, these results are difficult to interpret. Particularly, the utilization of GNNs remains to be validated. Similarly, Xu et al., 2022b demonstrated that their method was faster than the heuristics and close to optimal. However, they confirmed the benefits of GCNs, branch exchange, and action separation in an ablation study.
@@
Lastly, Cao et al., (2023) evaluated their physics-informed GAT-SAC on IEEE 33- and 119-node systems. Their approach outperforms other methods several control methods, including SAC variants and GCN-SAC in reducing voltage deviations and maintaining safe voltage levels, especially under noisy conditions. Ablation studies emphasized the importance of the GAT-based network and the added robustness from the deep autoencoder. Tests on the IEEE 119-node system confirmed the method’s scalability and effectiveness in larger networks.
5.1.2 Distribution Grid Control in Emergency Mode
@@
Another way to control grid voltages is load shedding, which involves deliberately disconnecting certain loads. This drastic measure is usually a last resort to avoid total blackouts. Conversely, in case of a full or partial blackout occurs, system restoration is needed to restart the grid. This involves coordinated steps to reconnect loads, restore generation capacity, and ensure the integrity of distribution networks.
@@
The goal is to develop a GRL agent for load shedding decisions, as outlined by Hossain et al., (2021), Zhang et al., (2023), and Pei et al., (2023), each proposing different GRL methods for managing critical system loads.
@@
The problem is framed as an MDP based on grid state observations like power demand, generation, voltage measurements, and topology. Pei et al., (2023) also use historical node voltages. Hossain et al., (2021) and Zhang et al., (2023) model decisions as binary (shed or not), while Pei et al., (2023) consider shedding 5% or 10% of the load. All approaches focus on heavy load nodes, with the action space scaling accordingly. Zhang et al., (2023) also include line switching.
The reward in all approaches is aimed at maintaining stable voltage levels and preventing system collapse. For example, Hossain et al., (2021) give a large negative reward if the voltage has not returned to nominal within a specified time, and a positive reward for voltages within predefined levels. The reward also minimizes load shedding or maximizes power supply (Zhang et al., (2023)) and penalizes actions that violate system constraints.
@@
The employed RL algorithms include DDQN with an $\epsilon$-greedy strategy and experience replay (Hossain et al., (2021) and Pei et al., (2023)), with Pei et al., (2023) also using a dueling architecture. Zhang et al., (2023) use PPO with a hybrid policy network combining a fully connected network and a GCN.
@@
All studies model the grid as a graph with nodes representing buses, including substations, loads, and generators connected by edges representing power lines and transformers. Node features include grid state data such as voltage measurements and loads. All approaches use a GCN for feature extraction but with different architectures. Hossain et al., (2021) use a simple GCN with three layers, while Zhang et al., (2023) employ a graph capsule network (cf. Sec. 3) to embed whole graphs and store additional neighborhood statistics, and Pei et al., (2023) GraphSAGE, which samples and aggregates the local neighborhood of nodes.
@@
Hossain et al., (2021) and Pei et al., (2023) use the IEEE 39-bus system, while Zhang et al., (2023) test on modified IEEE 13-bus and 34-bus systems. Each method trains on different topological configurations with random fault locations. All three approaches outperform feed-forward neural networks in convergence and average reward, especially on unseen topologies. Hossain et al., (2021) test on 32 topologies in the IEEE 39-bus system, while Pei et al., (2023) apply their GNN method to a 300-bus system, handling larger action spaces effectively. Zhang et al., (2023) and Pei et al., (2023) also outperform traditional optimization techniques in speed and near-optimal performance. Pei et al., (2023) show that GraphSAGE is more adaptable and efficient than classical GCN. These methods demonstrate improved voltage control across varying grid topologies, highlighting the robustness and efficiency of GNNs over traditional methods.
@@
On a different note, in case earlier mitigation actions fail and a blackout occurs, rapid system restoration is crucial to reconnect loads and restart the grid promptly. Zhao and Wang, (2022) employ a multi-agent approach where a Q-Network guides actions, using an encoder to observe generator capacities, switch statuses, and load conditions. A GCN extracts features from local and neighboring agents for the Q-Networks. This method outperforms single-agent DQN and other multi-agent baselines using feedforward networks, as well as CPLEX, a mathematical optimization method, in terms of accuracy and speed. Case studies validate this approach on IEEE 123 and 8500 node test systems Zhao and Wang, (2022).
@@ 5.2 Other Use Cases
Two additional use cases for distribution grids include loss minimization and economic dispatch. Loss minimization focuses on optimizing the power grid state by adjusting topology or generator set points to reduce losses in branch elements. Economic dispatch minimizes operational generator costs. Tab. 3 lists the RL method, action type, GNN architecture, grid size, and overall focus of the analyzed approaches
Jacob et al., (2022) address Distribution Network Reconfiguration (DNR) for loss minimization and resilience enhancement by re-configuring network topology via sectionalizing and tie line switches. The states include the network topology, connection status, power demands, generation outputs, bus voltages, branch currents, and total network loss. Rewards incentivize loss reduction, penalize disconnections or radial structure alterations, and reward feasible network exploration.
Meanwhile, Chen et al., (2023) focuses on economic dispatch in systems with high renewable energy, optimizing generation, renewable output, and ESS power under dynamic conditions. States include load demands, generation and RES outputs, ESS state of charge, and current timestep. Rewards optimize economic costs and system stability, addressing voltage violations and balancing economic operation costs and stability.
Graph representations in Jacob et al., (2022) denote substations and buses as nodes, while the edges represent power lines and transformers. They employ a Capsule-based Graph Convolutional Network (GCAPCN) (cf. Sec. 3) to capture local and global features. The GCAPCN encodes the features and topology for the policy network, while a feedforward neural network serves as the value network of a PPO algorithm. Likewise, Chen et al., (2023) represent the power grid the same way while using a GCN to process features like load demand, generation outputs, RES outputs, ESS state, and timestep information. Then, they apply the SAC algorithm with the GCN layers in both the actor and critic to optimize dispatch policies off-policy.
Table 3: Overview of other relevant GRL approaches for distribution grids. The RL algorithm, the action as well as the GNN type are specified. RES and ESS refer to renewable energy sources and energy storage system, respectively. The column focus defines the objective of the proposed approaches. DNR refers to Distribution Network Reconfiguration
Approach RL Action GNN Focus/ Unique Feature
@@
Jacob et al., (2022) evaluate their GCAPS-RL method against a feed-forward approach on modified IEEE 13- and 34-bus systems as well as two conventional baseline methods, Mixed-integer Second-order Conic Programming (MISOCP) and binary particle swarm optimization (BPSO). GCAPS-RL shows superior real-time decision-making and adherence to topological constraints and outperforms the feed-forward counterpart. Chen et al., (2023) conduct case studies on a modified IEEE39 system with conventional generators, renewable sources, and an ESS. Their GRL approach outperforms Optimal Solution with Perfect Information (OSPI), Model Predictive Control (MPC), and feed-forward SAC policies and shows strong convergence, effective policy performance, and superior scalability with cost reductions.
@@ 5.3 Discussion
Voltage and grid control in emergencies using DRL techniques involve several considerations. Reward functions typically address voltage deviations and may also integrate additional factors such as renewables or power loss. However, balancing these objectives poses a non-trivial challenge as mentioned in Sec. 4.5.
The control devices considered are mainly PV inverters, ESS, and generator power adjustments, with a notable approach that also includes topological actions. For zonal or microgrid distribution networks, multi-agent setups are very suitable as agents can represent different zones. Strategies vary in the handling of global system knowledge, using either global training with local evaluation or centralized critics. GNNs are used to integrate information from different agents as shown in Mu et al., (2023). The multi-agent approach matches the problem setup of zoned grids or microgrids very well and the conducted experiments show that GNNs improve the robustness of these methods.
GNNs implementations vary widely, with no consensus on their use in the actor, critic, or both. The node embeddings are mapped to action vectors through various readout methods, including neural networks, autoencoders, and 1D convolutions. Some approaches share weights between surrogate models and actors/critics. While graph representations are consistent, GNN architectures vary significantly, with spectral GCNs being common. Further research is therefore required to optimize the design of GNNs in voltage regulation use cases.
Temporal information is used in some approaches, but most only consider static data. The benefit of using temporal data remains an open question, as it typically increases the complexity of the models. So far, there are too few studies in this direction. Customized Graph Structure Operators, using domain knowledge, enhance feature extraction and define edge weights and critical connections. Customized GSOs are particularly promising as they leverage domain knowledge, enhancing feature extraction and graph representation for grids.
Similar to the conclusions in Sec. 4.5, GRL agents outperform DRL agents with fully connected neural networks in transferability and adaptability to topology changes, handling experiments with deleted grid elements or different structures without significant performance drops. GNNs’ ability to manage noisy or missing data is a major advantage, demonstrating robustness in experiments with generator failures, deleted lines, or nodes. This is particularly advantageous given the prevalence of faulty sensor data in real grid operations.
@@
The experiments and evaluations of these approaches cover a wide range of considerations. Most studies conduct experiments on IEEE grids, with grid sizes ranging from small systems with as few as 5 buses to large networks with up to 8500 nodes. There is a considerable variation in grid size, which affects the scalability and generalisability of the approaches. Notably, some methods show effective performance even without access to global information, highlighting the robustness of GRL strategies in this context. Evaluation metrics typically include voltage deviation, network energy loss, or voltage violation rates, reflecting the overarching goal of voltage control and load shedding. Traditional optimization techniques and heuristics serve as benchmarks in many studies, highlighting the comparative performance and efficiency gains of DRL-based approaches. In addition, comparisons with other DRL methods, including dense-based and CNN-based approaches, shed light on the relative advantages of graph-based methods. The experiments demonstrate the scalability, stability, and real-time capabilities of the proposed frameworks in addressing complex power system control challenges.
However, it should be noted that none of the studies have been carried out on real data. Most approaches rely on simplifications, such as considering only binary actions for load shedding. This clearly limits the applicability and highlights the need for experimentation in more realistic scenarios. Nevertheless, the studies presented confirm the potential of GRL for distribution system use cases.